# 3D Diffuser Actor: Multi-task 3D Robot Manipulation with Iterative Error Feedback

### v2

@@ A-E Detailed Model Diagram
The inputs to our network are i) a stream of RGB-D views; ii) a language instruction; iii) proprioception in the form of end-effector’s history poses; iv) the current noisy estimates of position and rotation; v) the denoising step $t$. The images are encoded into visual tokens using a pretrained 2D backbone. The depth values are used to “lift" the multi-view tokens into a 3D feature cloud. The language is encoded into feature tokens using a language backbone. The proprioception is represented as learnable tokens with known 3D locations in the scene. The noisy estimates are fed to linear layers that map them to high-dimensional vectors. The denoising step is fed to an MLP.
@@
The visual tokens cross-attend to the language tokens and get residually updated. The proprioception tokens attend to the visual tokens to contextualize with the scene information. We subsample a number of visual tokens using Farthest Point Sampling (FPS) in order to decrease the computational requirements. The sampled visual tokens, proprioception tokens and noisy position/rotation tokens attend to each other. We modulate the attention using adaptive layer normalization and FiLM [55]. Lastly, the contextualized noisy estimates are fed to MLP to predict the error terms as well as the end-effector’s state (open/close).
A-F Hyper-parameters for experiments
@@
PerAct GNFactor CALVIN Model $\mathrm{image\_size}$ $256$ $256$ $200$ $\mathrm{embedding\_dim}$ $120$ $120$ $192$ $\mathrm{camera\_views}$ $4$ $1$ $2$ $\mathrm{action\_history\_length}$ $3$ $3$ $3$ $\mathrm{FPS:\%\ of\ sampled\ tokens}$ $20\%$ $20\%$ $33\%$ $\mathrm{diffusion\_timestep}$ $100$ $100$ $25$ $\mathrm{noise\_scheduler:position}$ $\mathrm{scaled\_linear}$ $\mathrm{scaled\_linear}$ $\mathrm{scaled\_linear}$ $\mathrm{noise\_scheduler:rotation}$ $\mathrm{squaredcos}$ $\mathrm{squaredcos}$ $\mathrm{squaredcos}$ $\mathrm{action\_space}$ absolute pose absolute pose relative displacement Training $\mathrm{batch\_size}$ $240$ $240$ $1080$ $\mathrm{learning\_rate}$ $1e^{-4}$ $1e^{-4}$ $3e^{-4}$ $\mathrm{weight\_decay}$ $5e^{-4}$ $5e^{-4}$ $5e^{-3}$ $\mathrm{total\_epochs}$ $1.6e^{4}$ $8e^{5}$ $450$ $\mathrm{optimizer}$ Adam Adam Adam $\mathrm{loss\ weight}:w_{1}$ $30$ $30$ $30$ $\mathrm{loss\ weight}:w_{2}$ $10$ $10$ $10$ Evaluation $\mathrm{maximal\ \#\ of\ keyposes}$ $25$ $25$ $60$
@@
TABLE VI: Hyper-parameters of our experiments. We list the hyper-parameters used for training/evaluating our model on RLBench and CALVIN simulated benchmarks. On RLBench we conduct experiments under two setups: PerAct and GNFactor.
@@
Table VI summarizes the hyper-parameters used for training/evaluating our model. On CALVIN we observed that our model overfits the training data, resulting in lower test performance. We use higher $\mathrm{weight\_decay}$ and shorter $\mathrm{total\_epoch}$ on CALVIN compared to RLBench.
@@
A-G The importance of noise scheduler
We visualize the clean/noised 6D rotation representations as two three-dimensional unit-length vectors in Figure 7. We plot each vector as a point in the 3D space. We can observe that noised rotation vectors generated by the squared linear scheduler cover the space more completely than those by the scaled linear scheduler.

### v3

@@ B.3 Detailed Model Diagram of 3D Diffuser Actor
The inputs to our network are i) a stream of RGB-D views; ii) a language instruction; iii) proprioception in the form of current end-effector’s poses; iv) the current noisy estimates of position and rotation; v) the diffusion step $i$. The images are encoded into visual tokens using a pretrained 2D backbone. The depth values are used to “lift" the multi-view tokens into a 3D feature cloud. The language is encoded into feature tokens using a language backbone. The proprioception is represented as learnable tokens with known 3D locations in the scene. The noisy estimates are fed to linear layers that map them to high-dimensional vectors. The denoising step is fed to an MLP.
@@
The visual tokens cross-attend to the language tokens and get residually updated. The proprioception tokens attend to the visual tokens to contextualize with the scene information. We subsample a number of visual tokens using Farthest Point Sampling (FPS) in order to decrease the computational requirements. The sampled visual tokens, proprioception tokens and noisy position/rotation tokens attend to each other. We modulate the attention using adaptive layer normalization and FiLM [perez2018film]. Lastly, the contextualized noisy estimates are fed to MLP to predict the error terms as well as the end-effector’s state (open/close).
B.4 Detailed Model Diagram of Act3D
To make the paper self-contained, we present a detailed architecture diagram of Act3D in Figure 5. Act3D first encodes posed RGB-D images into 3D scene tokens and samples ghost points uniformly from a fixed-radius sphere, centered at the current end-effector position. Act3D initializes a learnable query and the ghost points with learnable latent embeddings, which are contexturalized by 3D scene tokens using 3D relative attention (Eqn. 1). Act3D measures the feature similarity of each ghost point and the learnable query to classify if a ghost point is close to the position of target end-effector pose, which is determined by the XYZ location of the ghost point with the highest score. Act3D then crops the scene with estimated end-effector position, and iterates the same classification procedure. The rotation and openess are predicted with learnable query embeddings.
B.5 Hyper-parameters for experiments
@@
Multi-view RLBench Single-view RLBench CALVIN Model    $\mathrm{image\_size}$ $256$ $256$ $200$    $\mathrm{embedding\_dim}$ $120$ $120$ $192$    $\mathrm{camera\_views}$ $4$ $1$ $2$    $\mathrm{FPS:\%\ of\ sampled\ tokens}$ $20\%$ $20\%$ $33\%$    $\mathrm{diffusion\_timestep}$ $100$ $100$ $25$    $\mathrm{noise\_scheduler:position}$ $\mathrm{scaled\_linear}$ $\mathrm{scaled\_linear}$ $\mathrm{scaled\_linear}$    $\mathrm{noise\_scheduler:rotation}$ $\mathrm{squaredcos}$ $\mathrm{squaredcos}$ $\mathrm{squaredcos}$    $\mathrm{action\_space}$ absolute pose absolute pose relative displacement Training    $\mathrm{batch\_size}$ $240$ $240$ $5400$    $\mathrm{learning\_rate}$ $1e^{-4}$ $1e^{-4}$ $3e^{-4}$    $\mathrm{weight\_decay}$ $5e^{-4}$ $5e^{-4}$ $5e^{-3}$    $\mathrm{total\_epochs}$ $1.6e^{4}$ $8e^{5}$ $90$    $\mathrm{optimizer}$ Adam Adam Adam    $\mathrm{loss\ weight}:w_{1}$ $30$ $30$ $30$    $\mathrm{loss\ weight}:w_{2}$ $10$ $10$ $10$ Evaluation    $\mathrm{maximal\ \#\ of\ keyposes}$ $25$ $25$ $60$
@@
Table 7: Hyper-parameters of our experiments. We list the hyper-parameters used for training/evaluating our model on RLBench and CALVIN simulated benchmarks. On RLBench we conduct experiments under multi-view and single-view setup.
@@
Table 7 summarizes the hyper-parameters used for training/evaluating our model. On CALVIN we observed that our model overfits the training data, resulting in lower test performance. We use higher $\mathrm{weight\_decay}$ and shorter $\mathrm{total\_epoch}$ on CALVIN compared to RLBench.
B.6 Denoising Diffusion Probabilistic Models
For self-completeness, we provide a formulation of denoising diffusion process. Please check [ho2020denoising] for more details. A diffusion model learns to model a probability distribution $p(x)$ by inverting a process that gradually adds noise to a sample $x$. For us, $x$ represents a sequence of 3D translations and 3D rotations for the robot’s end-effector. The diffusion process is associated with a variance schedule $\{\beta^{i}\in(0,1)\}_{i=1}^{N}$, which defines how much noise is added at each time step. The noisy version of sample $x$ at time $i$ can then be written as $x^{i}=\sqrt{\bar{\alpha}^{i}}x+\sqrt{1-\bar{\alpha}^{i}}\epsilon$ where $\epsilon\sim\mathcal{N}(\mathbf{0},\mathbf{1}),$ is a sample from a Gaussian distribution (with the same dimensionality as $x$), $\alpha^{i}=1-\beta^{i}$, and $\bar{\alpha}^{i}=\prod_{j=1}^{i}\alpha^{j}$. The denoising process is modeled by a neural network $\hat{\epsilon}=\epsilon_{\theta}(x^{i};i)$ that takes as input the noisy sample $x^{i}$ and the noise level $i$ and tries to predict the noise component $\epsilon$.
Diffusion models can be easily extended to draw samples from a distribution $p(x|\textbf{c})$ conditioned on input c, which is added as input to the network $\epsilon_{\theta}$. For us, c is the visual scene captured by one or more calibrated RGB-D images, a language instruction, as well as the proprioceptive information. Given a collection of $\mathcal{D}=\{(x^{i},\textbf{c}^{i})\}_{i=1}^{N}$ of end-effector trajectories $x^{i}$ paired with observation and robot proprioceptive context $\textbf{c}^{i}$, the denoising objective becomes:
$\mathcal{L}_{\text{diff}}(\theta;\mathcal{D})=\tfrac{1}{|\mathcal{D}|}\sum_{x^{i},\textbf{c}^{i}\in\mathcal{D}}||\epsilon_{\theta}(\sqrt{\bar{\alpha}^{i}}x^{i}+\sqrt{1-\bar{\alpha}^{i}}\epsilon,\textbf{c}^{i},i)-\epsilon||.$ (2)
This loss corresponds to a reweighted form of the variational lower bound for $\log p(x|\textbf{c})$ [DDPM].
In order to draw a sample from the learned distribution $p_{\theta}(x|\textbf{c})$, we start by drawing a sample $x^{N}\sim\mathcal{N}(\mathbf{0},\mathbf{1})$. Then, we progressively denoise the sample by iterated application of $\epsilon_{\theta}$ $N$ times according to a specified sampling schedule [DDPM, song2020denoising], which terminates with $x^{0}$ sampled from $p_{\theta}(x)$:
$x^{i-1}=\frac{1}{\sqrt{\alpha^{i}}}\left(x^{i}-\frac{\beta^{i}}{\sqrt{1-\bar{\alpha}^{i}}}\epsilon_{\theta}(x^{i},i,\textbf{c})\right)+\frac{1-\bar{\alpha}^{i+1}}{1-\bar{\alpha}^{i}}\beta^{i}\mathbf{z},$ (3)
where $\mathbf{z}\sim\mathcal{N}(\mathbf{0},\mathbf{1})$.
@@ B.6 Denoising Diffusion Probabilistic Models
B.7 The importance of noise scheduler
Existing diffusion policies [xian2023chaineddiffuser, chi2023diffusion] often deploy the same noise scheduler for estimate of position and rotation. However, we empirically found that selecting a good noise scheduler is critical. We visualize the clean/noised 6D rotation representations as two three-dimensional unit-length vectors in Figure 6. We plot each vector as a point in the 3D space. We show that noised rotation vectors generated by the squared linear scheduler cover the space more completely than those by the scaled linear scheduler, resulting in better performance than a square cosine scheduler